OpenStudy (anonymous):

|2x – 3| = 13? pleeeease

OpenStudy (zzr0ck3r):

|a| = b then a=b and a=-b so you have 2x-3=13 and 2x-3=-13 solve both for x x=8 and x=-5

OpenStudy (anonymous):

ohh thank you ! can u help me do this one ? P = 2w + 2h for w.

OpenStudy (zzr0ck3r):

subtract 2h from both sides, then divide by w on both sides

OpenStudy (anonymous):

how do u divide by w ?

OpenStudy (zzr0ck3r):

subtract 2h from both sides P-2h = 2w I meant divide by 2 (P-2h)/2=w

OpenStudy (anonymous):

w = P – 2h w = P – 2h – 2 w = the quantity P minus h, divided by two w = the quantity P minus two h, divided by two here are my choices

OpenStudy (zzr0ck3r):

the last one...

OpenStudy (anonymous):

thanks! can you show me how to solve this one step by step 2(3x – 4) = 3x + 7.